# Multi-scale Laplacian community detection in heterogeneous networks

### v3

@@ II Local modularity in heterogeneous networks
Since our new theoretical scheme is defined through a free parameter $\tau$ directly related to the scale of network structures, our method emerges as a natural tool for a systematic approach to this issue and the consequent detection of such local structures. This is done by introducing a Local Partition Stability Index $\Psi_{L}(\tau)$ for each local cluster in the dendrogram at time $\tau$. It is defined as the length in terms of $\log\Delta$ of the branch of the dendrogram between the creation of a cluster by the merging of subclusters and the further merging of such cluster in a supercluster (see Fig. 4). It is essential to underline that, as explained above, $\Psi_{L}(\tau)$ substantially measures cluster the Shannon information contained by a cluster related to distribution of distances induced by diffusion evolution operator $e^{-\tau\hat{L}}$ in the cluster. In other words, the higher the value of $\Psi_{L}(\tau)$, the more connected by diffusion is such cluster on the time-scale $\tau$. Figure 4(a) shows a particular example of Zachary’s karate club using the Laplacian, $\hat{L}$. As we illustrate in Fig. 4(b), this module exhibits a large $\Psi_{L}(\tau)$ for all the expected range of significative timescales $\tau$ both for $\hat{L}$ and $\hat{L}_{RW}$ (see Sect. V below).
@@ III Temporal dendrograms and metastable nodes
In this section, we face the problem of the stability of the best cut of the dendrograms at varying diffusion time $\tau$. As previously discussed, one central implication of the LRG framework is its dynamic perspective on the contribution of single nodes to diffusionally connected mesoscopic structures, when the time-scales $\tau$ is changed. In particular, by increasing $\tau$, a single node may always contribute to a single growing diffusionally connected structure or cluster, or it can suddenly switch cluster at a certain $\tau$. We focus on this second crucial kind of nodes that, precisely for this property, acts as a communication bridge between stably growing network modules. For this reason, we call this subset of nodes as metastable nodes. They permit different highly connected subnetworks to share information. Consequently, metastable nodes contribute to different functional communities at different diffusional resolution scales $\tau$ of the network structures. This explains why classifying these nodes is exceptionally challenging when using standard optimization algorithms that provide a static picture of the network. Due to the time-dependent definition of the distance matrix $\mathcal{D}_{ij}(\tau)$, our method is instead very effective in identifying this class of metastable nodes through the direct analysis of the composition of the clusters corresponding to the best partition of the dendrogram at increasing $\tau$.
@@
For instance, the Sankey diagram in Fig. 5 illustrates this phenomenon well, showing that Node 3 of Zachary’s Karate club at a specific $\tau$ changes module it belongs to. The figure clearly shows that the best community assignment for this particular node changes at a certain time $\tau$ between $\tau^{\prime}=1/\lambda_{max}$ and the peak of the specific heat at $\tau^{*}$, revealing its role as “bridge” node. This feature assigns it a central role in communication and/or control processes between the two weakly dependent modules. In more general cases, this makes it plausible to expect a similar relevance of metastable nodes in real networks (e.g., the human connectome) for synchronizability and information integration. These properties of the just defined metastable nodes recall and generalize the concept of “modular flexibility” as defined for the dynamic correlation analysis of time-series in [55] or previously described in community degeneracy problems [56, 57]. The modular flexibility of a given node in the correlation network represents how frequently it changes the module it belongs to across time. It means that nodes more likely to be connected to multiple modules at different times are more flexible. Modular flexibility has been associated with network adaptability, giving flexible nodes the role of pivoting the dynamics processes running on top of the network.
@@
IV Identifying communities in real-world networks
In this section, first we summarize the fundamental steps of our new framework for community detection. Then, we illustrate its application to two benchmark protein-protein interaction (PPI) networks to detect their global and local intrinsic modules. The practical application of the MSL framework can be summarized in the following fundamental steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
As important examples of real networks, we apply this procedure to the complex hierarchical structure of the E. Coli and M. Musculus PPI networks [58], whose results are shown in Figure 6. We identify the functional role of these communities through the use of the UniProt Knowledgebase [59], a comprehensive, high-quality, and freely accessible set of around 190 million protein sequences annotated with functional information. In particular, we use the keywords associated with each protein to facilitate the characterization of different core functionalities of the observed modules. The PPI network of E.Coli presents a highly nested tree-like structure with heterogeneous submodules dedicated to different biological functions, as shown in Figure 6(a). For clarity, we explore the division into five modules exemplified in the dendrogram of Figure 6(a). In this case, two main communities emerge: the first one related to RNA-binding (green nodes,  88% of the proteins engaged in this process belong to this community, constituting  59% of it), and the second one inherent to chemotaxis and DNA repair (violet nodes, containing the 80% of the proteins involved in chemotaxis, and conforming the 82% of the community). We also identify two different particular metastable nodes (see black nodes in Fig. 6(a)): ’P0AGB6’ a $\sigma^{E}$-factor responsible for the envelope stress response and ’P0A6M8’, an Elongation factor G that catalyzes the GTP-dependent ribosomal translocation step during translation elongation. On the contrary, the network of M.Musculus appear much more interconnected and highly intricate. However, we can still discern three principal modules, the smallest one of which (blue nodes in Figure 6(b)) is strongly related to the signal processing function and the immune response of the animal’s cells. Finally, the direct application of the LPSI $\Psi_{L}(\tau)$ at the different branches of the dendrogram allows the detection of a robust, independent local module (see black dashed lines in Fig. 6(b)), composed of a group of proteins dedicated to the Electron Transport Chain, a driver of the adenosine triphosphate (ATP) synthesis in mithocondria of eukaryotic cells. Even though a detailed study of different biological networks is beyond the scope of this manuscript, these results appear to be highly illustrative of the power of the new proposed method. We will leave more refined analyses of these nested structures to single important cases to future works.
V The Laplacian Random-Walk
@@
In the previous sections, we have illustrated our new method of community detection based on the time-dependent metric induced by the graph diffusion evolution operator $e^{-\tau\hat{L}}$. One may wonder if this method can be extended to other physically meaningful dynamics of exploration of the network. This is the case, for instance, for the stochastic evolution of random walks, which is described by the random walk Laplacian $\hat{L}_{RW}=\hat{D}^{-1}\hat{L}$. This encodes a discrete time-traveling dynamics of a walker moving from a node to any nearest neighbor node with uniform probability: i.e., the transition matrix for the random walk dynamics on top of a graph [60, 61] (see Appendix E for more details). It is important to note that the operator $\hat{\rho}(\tau)$ of Eq. (1), on which the above method is based, is properly defined because $\hat{L}$ is symmetric and positive semi-definite, while $\hat{L}_{RW}$ in general is not. Even though it is not symmetric, all eigenvalues of $\hat{L}_{RW}$ are real and satisfy $0\leq\mu_{i}\leq 2$ with $\mu_{i}=2$ included only in the case of bipartite networks and $\hat{L}_{RW}$ is similar to the symmetric operator $\hat{L}_{sym}=D^{1/2}\hat{L}_{RW}\hat{D}^{-1/2}$. In this way, we can reformulate our method for the random walk exploration dynamics of the network by simply substituting $\hat{L}$ with $\hat{L}_{sym}$.
@@
When the spectral density $p(\mu)$ of $\hat{L}_{RW}$ (or equivalently of $\hat{L}_{sym}$) is power-law distributed at small values of $\mu$, the return-time distribution of the random walker, i.e., the probability $P_{0}(t)$ that the first return happens after $t$ time steps, allows the definition of the random walk spectral dimension ($d_{S}$) of a network [62, 63] from $P_{0}(t)\propto t^{-d_{S}/2}$, as a natural generalization of what happens in a $d-$dimensional lattice for which $P_{0}(t)\propto t^{-d/2}$ [63]. In general, heterogeneous networks with sufficiently high connectivity are characterized by a spectral density that takes the well-known semicircular shape (Wigner semicircle law) in the interval $[2-\mu_{max},\mu_{max}]$, which is centered at $\mu=1$ [64, 65, 66, 67], when $d_{S}$ is large enough. Consequently, to better analyze and compare the spectrum of $\hat{L}_{RW}$ of different real networks, it is convenient to standardize the eigenvalues so that the semicircular random parts of the spectrum, if present, overlap in all analyzed cases. This is expected in the case of networks with a large enough lowest degree [68] (roughly corresponding to a sufficiently large effective value of $d_{S}$). This standardization consists of the simple transformation $\mu^{\prime}=\frac{1}{2}+\frac{\mu-1}{2(\mu_{\max}-1)}$. By contrast, the negative part of the spectrum in the standardized eigenvalues $\mu^{\prime}$ (with the exception of the trivial lowest eigenvalue $\mu=0$ which corresponds to $\mu^{\prime}=\frac{\mu_{max}-2}{2(\mu_{\max}-1)}<0$) is absent for completely random cases, while it is expected when rare regions effects –often leading to anomalous slow relaxation of the system– are present [69] (see also Appendix E). We illustrate in Fig. 7(d) different benchmark monopartite networks where the spectrum of $\hat{L}_{RW}$ converges to this expected universal shape  [^2] Indeed, as the density matrix encodes pairwise information exchange between nodes, we expect the existence of such universal behavior for an effective spectral dimension $d_{S}^{*}=4$, similarly to the Gaussian model [74, 32, 63, 76].
@@
Figure 7(a) shows the “random-walk” dendrogram for Zachary’s karate club by considering $\hat{L}_{RW}$. It is important to note that, at a time equal to the inverse of the maximum eigenvalue, the Node 3 is placed in a different community in the two cases of using $\hat{L}$ or $\hat{L}_{sym}$ in the metrics (compare with Fig. 2(c)) indirectly confirming its metastable nature. Figure 7(b) shows the entropy $S_{RW}(\tau)$ calculated now using the random walk density matrix $\hat{\rho}_{RW}(\tau)=\exp(-\tau\hat{L}_{sym})/Z_{RW}(\tau)$, and the related susceptibility (i.e. specific heat) $C_{RW}(\tau)$ for the same network. Note that $C_{RW}$ presents a slightly different shape concerning the one of Fig. 2(b) obtained by the Laplacian $\hat{L}$, as an effect of the dynamical modes of the network induced by $\hat{L}_{sym}$ (i.e. $\hat{L}_{RW}$). Figure 7(c) compares the particular time $\tau^{\prime}=1/\mu_{max}$ as determined by the maximization of $\Psi(n;\tau)$ (see Fig. 7(e) for an analysis of the two first optimal gaps over multiple timescales). Figure 4(b) shows the comparison of the behavior of the Local Partition Stability Index $\Psi_{L}(\tau)$ for the same branch of the dendrograms at different times using both $\hat{L}$ and $\hat{L}_{sym}$. Besides, we have also characterized the characteristic time of change for the metastable Node 3, $\tau=1.94(1)\simeq 3.31\tau^{\prime}$.
@@
Figures 7(f) and (g) show the comparison of the two Laplacian entropies and the related specific heats (or entropic susceptibilities) for the Mus Musculus PPI network [58] and for the color vision circuit in the medulla of Drosophila Melanogaster [71]. They show how information is differently integrated in the two networks by the two Laplacians. In particular, these effects will strongly depend on the effective (spectral) dimension of the network, both at global or local scales [72]. As a simple and illustrative example of this effect, one can consider the specific analysis of the Star-Clique network (see Appendix F), where a clique of $m\gg 1$ nodes is connected to a star of $n\gg 1$ nodes by a single bridge edge: a random walker will be effectively trapped at the star hub or in the clique for a long time before visiting both structures. This leads to two well-defined and dynamically separated communities using $\hat{L}_{RW}$. Instead, the diffusion flow with a uniform rate per edge, described by the “continuous” Laplacian $\hat{L}$, will additionally capture the broadcast of information linked to both hubs, giving place to more accurate information on the topological properties of the network (see further details in Appendix F).
In general, we can say that the major difference is that, while $\hat{L}$ is extensive in the local connectivities of the network, $\hat{L}_{RW}$ is intensive and the (stochastic) flow is normalized over the whole neighborhood of each single node. This gives rise to trapping mechanisms in the random walk when two highly connected subregions only communicate weakly.
@@
Another important aspect that must be considered when deciding which of the two operators is more appropriate to be used is their relation to eventual dynamical models on the network to be studied. Indeed, in general, statistical and dynamical models on graphs, such as epidemic and contact processes, Ising and Kuramoto models, when written in terms of continuous fields on the nodes of the graph present a two-body interaction term, giving the Gaussian approximation, substantially determined by $\hat{L}$. We emphasize that this is consistent with previous studies underlining the relevance of using $\hat{L}$ instead of Ł$\L_{RW}$ for community detection purposes [26]. This is because if the strength of the interaction between any pairs of neighboring nodes, is fixed, then the global interaction that a single node has with the rest of the network is extensive in the local connectivity, i.e. the node degree. This is the case of the Laplacian $\hat{L}$, while in the operator $\hat{L}_{RW}$ it is the global interaction that any single node has with the rest of the network to be fixed to a unitary constant. For this reason, when one is interested in the best network partition for this kind of systems, the natural choice is to use $\hat{L}$ itself.
VI Discussion and conclusion
@@
The strong topological heterogeneity constitutes a fundamental issue in real networks [73], inducing non-unique and highly complex algorithms for partitioning and community detection. Grounding on the intrinsic property of the Laplacian time-evolution operator $e^{-\tau\hat{L}}$ to be a sort of telescopic ”scanner” of the mesoscopic scales of a network, we propose a simple but efficient algorithm based on the Laplacian Renormalization Group theory [33, 17], which permits the introduction of an ultrametric communication distance related to the diffusion of information to tackle the community detection problem. It is important to note that the same theoretical framework allows us to formulate a natural and parsimonious interpretation of the so-called resolution limit [24, 65] that generally affects the modularity community detection algorithm: it has been shown that using the modularity function is equivalent to limiting the weighting scheme for the diffusional paths to the linear contribution in the Laplacian operator. On the contrary, weighting the different paths from vertex $i$ to vertex $j$ through the operator $e^{-\tau\hat{L}}$ implies properly considering the contributions coming from all the significant powers of $\hat{L}$ for each value of the scale parameter $\tau$. In this way, by varying $\tau$ from the minimal significant value $\tau^{\prime}=1/\lambda_{max}$ and the maximal one $\tau^{*}$, corresponding to the large time peak of $C(\tau)$, we can resolve all the telescopic community structure of the network. This process also explains the defocusing of small-scale structures by increasing $\tau$: once a certain number of diffusive eigenmodes has been integrated out, it results in a loss of information about the microscopic network scales related to that part of the spectrum.
@@
We highlight important theoretical implications of our new approach. First of all, the LRG framework sets the substantial, even though not rigorous, equivalence between real-space methods that can be derived from the $\hat{\rho}(\tau)$-density matrix (i.e., Louvain methods for community detection or Markov stability) and spectral clustering techniques [18, 19, 20], which addresses the issue in $k-space$ and analyzes diffusion modes of the network. These approaches can be seen as two sides of the same coin. On the other hand, it allows proposing a unified interpretation of community detection in terms of structural and dynamical aspects, providing an overall vision over different methodologies that have been hitherto considered separately. We underline that the interaction matrix of the Gaussian field theory on graphs [74, 75] when the global interaction that a single node has with the rest of the network is extensive in the local connectivity, as the Gaussian approximation of any statistical model on graphs (Ising, Kuramoto, contact-epidemic models, etc), is always defined in terms of the Laplacian $\hat{L}$ As a natural consequence, we stress the general importance of using $\hat{L}$ –and not $\hat{L}_{RW}$– for community detection purposes on top of matrices resulting from dynamical processes in which the interaction of the single nodes with the rest of the networks is the sum of the uniform interactions with all neighbors. The use of $\hat{L}_{RW}$ has to be limited to the random walk itself and related stochastic models for the diffusion of single particles on the network substrate.
@@
In addition, it is also important to note that in general the spectrum of eigenvalues of $\hat{L}_{RW}$ takes the universal semicircular shape [64], independently of the peculiar topology of the network, when the minimal connectivity is sufficiently large [68] (it has been shown that this quantity controls the emergent functional communities and dynamical aspects of the system [63]). When the mean and minimal connectivity are finite and sufficiently large, the spectrum does not give local topological information on the heterogeneous network. This is because when the local connectivity is sufficiently large, a single random walk will visit just a negligible fraction of the network in the large time limit. In other words, different realizations of the process will give completely different information on the local topology of the network since their intersection is negligible. This point has been discussed in detail in [76] in the context of Euclidean spaces, where a strict relation with the upper critical dimension of $\phi^{4}$ field theory is built. This topological limitation does not affect the Laplacian $\hat{L}$ which is always sensitive to the local topological features. The reason grounds on its extensivity in the local connectivity of nodes. Indeed, independently of the minimal connectivity of the heterogeneous networks the fluid Laplacian $\hat{L}$ [22], through its time-evolution operator $e^{\tau\hat{L}}$, describes a process that progressively invades all the network.
@@
A further crucial feature to consider about the structural organization of heterogeneous networks is the possible existence of a subset of metastable nodes that, depending on the scale of observation of the network, can switch the community which they belong to. In principle, understanding the general features of this class of nodes can be particularly relevant for controllability aims in complex networks. For instance, one can expect this switching behavior on different scales $\tau$ to play a central role in forming attractors as expected for synchronization dynamics in hierarchic modular networks [77, 78]. A similarly central role may be played in the metastable dynamics in human brain networks, where hubs have been elucidated to control the resulting wave patterns [79]. We conjecture that this subset of nodes may be involved in the observed disrupted hub organization in the topological disturbances associated with schizophrenia [80], where the waste of hubs [81] can lead to dysfunctions in controlling and integrating neuronal signals [82]. Finally, our new multi-scale Laplacian (MSL) approach sheds light on the origin of the resolution limit problem while giving a global overview of the importance of local communities in the entire network [24].

### v4

@@ II Local modularity in heterogeneous networks
Since our new theoretical framework depends on a free parameter $\tau$ directly related to the scale of internal structures, it provides a natural solution for a systematic approach to this issue. We can, thus, introduce the Local Partition Stability Index $\Psi_{L}(\tau)$ for each local cluster in the dendrogram at time $\tau$. It is defined as the length in terms of $\log\Delta$ of the branch of the dendrogram between the creation of a cluster by the merging of subclusters and the further merging of such cluster in a supercluster (see Fig. 4). We must underline that, as explained above, $\Psi_{L}(\tau)$ measures the Shannon information contained by a cluster related to the distribution of distances induced by diffusion evolution operator $e^{-\tau\hat{L}}$ in the cluster. In short, the higher the value of $\Psi_{L}(\tau)$, the more connected by diffusion is such cluster on the time-scale $\tau$. Figure 4(a) shows a particular example of Zachary’s karate club using the Laplacian, $\hat{L}$. As we illustrate in Fig. 4(b), this module exhibits a large $\Psi_{L}(\tau)$ for all the range of significative timescales $\tau$ both for $\hat{L}$ and $\hat{L}_{RW}$ (see below).
@@ III Temporal dendrograms and metastable nodes
As previously discussed, one central implication of the LRG framework is its dynamic perspective on the contribution of single nodes to diffusionally connected mesoscopic structures when the time-scales $\tau$ change. Mesoscopic divisions roughly maintain their internal composition (i.e., they are stable), but specific nodes can suddenly switch clusters at a certain $\tau$. We focus on such nodes that, precisely for this property, act as a communication bridge between network regions. Therefore, we denominate this subset of nodes as metastable nodes, as they permit different highly connected subregions to share information. Consequently, metastable nodes can be attributed to different subdivisions at different resolution scales of the system. This explains why classifying these nodes is exceptionally challenging when using standard optimization algorithms that provide a static picture of the network. Due to the time-dependent definition of the distance matrix $\mathcal{D}_{ij}(\tau)$, it is straightforward to identify metastable nodes through the direct analysis of the cluster composition corresponding to the best partition of the dendrogram at increasing $\tau$.
@@
The Sankey diagram in Fig. 5 illustrates this phenomenon, showing how Node 3 in Zachary’s Karate club changes the module it belongs to at a specific $\tau$. The figure clearly shows how the best assignment for this particular node changes at a specific time $\tau$ between $\tau^{\prime}=1/\lambda_{max}$ and the peak of the specific heat at $\tau^{*}$, revealing its role as “bridge” node. This feature assigns it a central role in communication and control processes between the two weakly dependent modules. In other cases, it is plausible to expect a similar relevance of metastable nodes in real networks (e.g., the human connectome) for synchronizability and information integration. The concept of metastable nodes generalize the one of “modular flexibility” as defined for the dynamic correlation analysis of time-series in [55] or previously described in community degeneracy problems [56, 57]. The modular flexibility of a given node in the correlation network represents how frequently it changes the module it belongs to across time. This means that nodes are more likely to be connected to multiple modules at different times and are more flexible, being associated with network adaptability, and giving flexible nodes the role of pivoting the dynamics processes running on top of the network.
@@
Disentangling PPI networks
As important examples of real networks, we apply the full methodology described above (summarized in Appendix D) to the complex structure of E. Coli and M. Musculus PPI networks [58], as shown in Figure 6. We identify the functional role of these communities via the UniProt Knowledgebase [59], a comprehensive, high-quality, and freely accessible set of around 190 million protein sequences annotated with functional information. In particular, we use the keywords associated with each protein to facilitate the characterization of different core functionalities of the observed modules. The PPI network of E.Coli presents a highly nested tree-like structure with heterogeneous submodules dedicated to other biological functions, as shown in Figure 6(a). For clarity, we explore the division into five modules exemplified in the dendrogram of Figure 6(a). In this case, two central communities emerge. The first one is related to RNA-binding (green nodes,  88% of the proteins engaged in this process belong to this community, constituting  59% of it), and the second one is inherent to chemotaxis and DNA repair (violet nodes, containing the 80% of the proteins involved in chemotaxis, and conforming the 82% of the community). We also identify two different particular metastable nodes (see black nodes in Fig. 6(a)): ’P0AGB6’, a $\sigma^{E}$-factor responsible for the envelope stress response, and ’P0A6M8’, an elongation factor G that catalyzes the GTP-dependent ribosomal translocation step during translation elongation. Conversely, the M.Musculus network appears much more interconnected and highly intricate. Still, we can discern three principal modules, the smallest one of which (blue nodes in Figure 6(b)) is strongly related to the signal processing function and the immune response of the animal’s cells. Finally, the direct application of $\Psi_{L}(\tau)$ allows the detection of a robust, independent local module (see black dashed lines in Fig. 6(b)), composed of a group of proteins dedicated to the Electron Transport Chain, a driver of the adenosine triphosphate (ATP) synthesis in mitochondria of eukaryotic cells. Even though a detailed study of different biological networks is beyond the scope of this manuscript, our results are highly illustrative of the power of the LRG, leaving extensive analyses of real structures for future works.
IV The Laplacian Random-Walk
@@
One might ask whether the current use of the LRG can be extended to other meaningful dynamics on the network. To do so, we examine the stochastic behavior of random walks, using the random walk Laplacian $\hat{L}_{RW}=\hat{D}^{-1}\hat{L}$. This encodes a discrete time-traveling dynamics of a walker moving from a node to any nearest neighbor with uniform probability: i.e., the transition matrix for a RW on top of a graph [60, 61] (see also Appendix F). The operator $\hat{\rho}(\tau)$ is well-defined because, although $\hat{L}_{RW}$ is not symmetric, all its eigenvalues are real and fall within the range $0\leq\mu_{i}\leq 2$. Notably, $\hat{L}_{RW}$ is similar to the symmetric operator $\hat{L}_{sym}=D^{1/2}\hat{L}_{RW}\hat{D}^{-1/2}$, enabling us to reformulate our framework for the RW exploration dynamics of the network by substituting $\hat{L}$ with $\hat{L}_{sym}$.
@@
In general, heterogeneous networks with sufficiently high connectivity are characterized by a spectral density that takes the well-known semicircular shape (Wigner semicircle law) in the interval $[2-\mu_{max},\mu_{max}]$, which is centered at $\mu=1$ [62, 63, 64, 65]. This suggests a common universal dynamic among these diverse structures. To better analyze the spectrum of $\hat{L}_{RW}$, the eigenvalues spectrum can be standardized to collapse the semicircular (Gaussian) part of the spectrum. A perfect collapse is expected for networks with a large enough lowest degree [66] (roughly corresponding to a sufficiently large effective value of the spectral dimension, $d_{S}$, see Appendix F). The standardization consists of the simple transformation $\mu^{\prime}=\frac{1}{2}+\frac{\mu-1}{2(\mu_{\max}-1)}$. Instead, the negative part of the spectrum in the standardized eigenvalues $\mu^{\prime}$ (except the trivial lowest eigenvalue $\mu=0$ corresponding to $\mu^{\prime}=\frac{\mu_{max}-2}{2(\mu_{\max}-1)}<0$) becomes absent for completely random cases, while it is expected when rare regions effects –often leading to anomalous slow relaxation of the system– are present [67] (see Appendix F). Fig. 7(d) reports different benchmark monopartite networks where the spectrum of $\hat{L}_{RW}$ converges to this expected universal shape  [^2] Indeed, as the density matrix encodes pairwise information exchange between nodes, we expect the existence of such universal behavior for an effective spectral dimension $d_{S}^{*}=4$, similarly to the Gaussian model [72, 32, 74, 75].
@@
Figure 7(a) shows the RW exploration dendrogram for Zachary’s karate club. We highlight that, for $\tau$ equal to the inverse of the maximum eigenvalue, Node 3 belongs to a different community using $\hat{L}$ or $\hat{L}_{sym}$ (comparing with Fig. 2(c)), indirectly confirming its metastable nature. Figure 7(b) shows the entropy $S_{RW}(\tau)$ calculated now using the RW density matrix $\hat{\rho}_{RW}(\tau)=\exp(-\tau\hat{L}_{sym})/Z_{RW}(\tau)$, and the related susceptibility (i.e. specific heat) $C_{RW}(\tau)$ for the same network. Note that $C_{RW}$ presents a slightly different shape compared to the one of Fig. 2(b) obtained by the Laplacian $\hat{L}$, as an effect of the dynamical modes of the network induced by $\hat{L}_{sym}$ (i.e. $\hat{L}_{RW}$). Figure 7(c) compares the particular time $\tau^{\prime}=1/\mu_{max}$ determined by the maximization of $\Psi(n;\tau)$ (see Fig. 7(e) for an analysis of the first two optimal gaps over multiple timescales). Figure 4(b) shows the comparison of the behavior of the Local Partition Stability Index $\Psi_{L}(\tau)$ for the same branch of the dendrograms at different times using both $\hat{L}$ and $\hat{L}_{sym}$. In addition, we have also characterized the characteristic change time for metastable Node 3, $\tau=1.94(1)\simeq 3.31\tau^{\prime}$.
@@
Figures 7(f) and (g) show the comparison of the two Laplacian entropies and specific heats (or entropic susceptibilities) for the Mus Musculus PPI network [58] and the color vision circuit in the medulla of Drosophila Melanogaster [69]. They show how the two Laplacians integrate information differently in the two networks. In particular, these effects will strongly depend on the effective (spectral) dimension of the network, both at global or local scales [70]. As a simple and illustrative example of this effect, one can consider the analysis of the Star-Clique network (see Appendix G): a clique of $m\gg 1$ nodes is connected to a star of $n\gg 1$ nodes by a single bridge edge. A random walker will be effectively trapped at the star hub or in the clique for a long time before visiting both structures. This leads to two well-defined and dynamically separated communities using $\hat{L}_{RW}$. Instead, the diffusion flow with a uniform rate per edge, described by the Laplacian $\hat{L}$, which additionally captures the broadcast of information linked to both hubs, giving place to more accurate information on the topological properties of the network (see Appendix G).
@@
The major difference between $\hat{L}$ and $\hat{L}_{RW}$ is that $\hat{L}$ is extensive in the local connectivities of the network while $\hat{L}_{RW}$ is intensive, and the stochastic flow is normalized over the entire neighborhood of every single node. This gives rise to trapping mechanisms in the RW when two highly connected subregions communicate weakly. When choosing which of the two operators is more appropriate, another aspect that must be considered is their relation to eventual dynamical models on the network to be studied. Statistical and dynamical models on graphs, such as epidemic and contact processes, Ising and Kuramoto models, when written in terms of continuous fields on the nodes of the graph, present a two-body interaction term, giving the Gaussian approximation, substantially determined by $\hat{L}$. This is consistent with previous studies underlining the relevance of using $\hat{L}$ instead of Ł$\L_{RW}$ for community detection purposes [26]. This is because if the strength of the interaction between any pairs of neighboring nodes is fixed, then the global interaction that a single node has with the rest of the network is extensive in the local connectivity, i.e., the node degree. This is the case of the Laplacian $\hat{L}$, while in the operator $\hat{L}_{RW}$, it is the global interaction that any single node has with the rest of the network to be fixed to a unitary constant. For this reason, when one is interested in the best network partition for this kind of systems, the natural choice is to use $\hat{L}$.
V Discussion and conclusion
@@
The strong topological heterogeneity constitutes a fundamental issue in real networks [71], inducing non-unique and highly complex algorithms for partitioning and community detection. Grounding on the intrinsic property of the Laplacian time-evolution operator $e^{-\tau\hat{L}}$ to be a sort of telescopic ”scanner” of the mesoscopic scales of a network, we propose a simple but efficient application of the Laplacian Renormalization Group theory [33, 17], which permits the introduction of an ultrametric communication distance related to the diffusion of information to disentangle the mesoscopic structure of any heterogeneous system. This naturally offers a natural and parsimonious interpretation of the so-called resolution limit [24, 63] that generally affects modularity community detection algorithms: it has been shown that using the modularity function is equivalent to limiting the weighting scheme for the diffusional paths to the linear contribution in the Laplacian operator. On the contrary, weighting the different paths from vertex $i$ to vertex $j$ through the operator $e^{-\tau\hat{L}}$ implies properly considering the contributions coming from all the significant powers of $\hat{L}$ for each value of the scale parameter $\tau$. Hence, by varying $\tau$ from the minimal significant value $\tau^{\prime}=1/\lambda_{max}$ and the maximal one $\tau^{*}$, corresponding to the large time peak of $C(\tau)$, we can resolve all the telescopic community structure of the network. This explains the defocusing of small-scale structures by increasing $\tau$. Once a certain number of diffusive eigenmodes has been integrated out, it results in a loss of information about the microscopic network scales related to that part of the spectrum.
Our framework has important theoretical implications. Firstly, it establishes a substantial (but not rigorous) equivalence between real-space methods derived from the $\hat{\rho}(\tau)$-density matrix (i.e., Louvain methods for community detection or Markov stability) and spectral clustering techniques [18, 19, 20], which addresses the issue in $k-space$ and analyzes diffusion modes of the network. These methods can be seen as two sides of the same coin. Furthermore, it provides a unified interpretation of community detection in terms of structural and dynamic aspects, offering an overall vision of different methodologies that were previously considered separately.
@@
We underline that the interaction matrix of the Gaussian field theory on graphs [72, 73] is always defined in terms of the Laplacian $\hat{L}$ when the global interaction that a single node has with the rest of the network is extensive in the local connectivity. This is the basis of the Gaussian approximation of any statistical model on graphs (Ising, Kuramoto, contact-epidemic models, etc). As a result, we stress the general importance of using $\hat{L}$ (not $\hat{L}_{RW}$) for community detection purposes on top of matrices resulting from dynamical processes in which the interaction of the single nodes with the rest of the networks is the sum of the uniform interactions with all neighbors. The use of $\hat{L}_{RW}$ has to be limited to the RW and related stochastic models for the diffusion of single particles on the network substrate.
@@
As a direct consequence, the eigenvalues spectrum of $\hat{L}_{RW}$ takes the universal semicircular shape [62], independently of the peculiar topology of the network, when the minimal connectivity is sufficiently large [66]. This quantity has been shown to control the emergent functional communities and dynamical aspects of the system [74]. We emphasize that the spectrum does not provide any local topological information on the heterogeneous network, as long as the mean and minimal connectivity are finite and sufficiently large: a single RW will visit a negligible fraction of the network in the large time limit. This means that different realizations of the process may yield completely different information on the local topology of the network, since their intersection is negligible. A similar point has been addressed in [75], in the context of Euclidean spaces, where a strict relation with the upper critical dimension of $\phi^{4}$ field theory is established. This topological limitation does not affect the Laplacian $\hat{L}$, which is always sensitive to local topological features. The reason grounds on its extensivity in the local connectivity of nodes. Indeed, independently of the minimal connectivity of the heterogeneous networks, the fluid Laplacian $\hat{L}$ [22], through its time-evolution operator $e^{\tau\hat{L}}$, describes a process that gradually invades all the network.
@@
We have evidenced a further crucial feature of heterogeneous networks: the existence of a subset of metastable nodes that, depending on the scale of observation of the network, can switch the community to which they belong. Understanding the general features of this class of nodes can be particularly relevant for controllability aims. For instance, one can expect this switching behavior on different scales $\tau$ to play a central role in forming attractors as expected for synchronization dynamics in hierarchic modular networks [76, 77]. A similar central role may be played in the metastable dynamics in human brain networks, where hubs have been elucidated to control the resulting wave patterns [78]. We conjecture that this subset of nodes may be involved in the observed disrupted hub organization in the topological disturbances associated with schizophrenia [79], where the waste of hubs [80] can lead to dysfunctions in controlling and integrating neuronal signals [81].